# Beyond Tools: Understanding How Heavy Users Integrate LLMs into Everyday Tasks and Decision-Making

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Recent studies have attempted to measure the impact of LLM use through quantitative metrics such as task performance and decision accuracy (Kim et al., 2025). However, these measurement-based approaches cannot fully capture how people delegate everday decisions to LLMs or the resulting meta-cognitive effects. Furthermore, everyday decisions encompass a broad spectrum of choices, from routine task management to social interaction planning (Eigner and Händler, 2024; Dhami and Thomson, 2012), making them challenging to examine through purely quantitative and task-specific approaches.
@@
  • RQ2: What underlying needs do LLM-heavy users seek to fulfill through LLM assistance?
  • RQ3: How do users conceptualize and evaluate their relationship with LLMs?
@@ 3.1.3. Case 3: Interpersonal Guidance through LLMs
3.2. RQ2: What underlying needs do LLM-heavy users seek to fulfill through LLM assistance?
@@ 3.2.1. Gaining Self-Confidence through Validation
Participants often consulted LLMs not to make decisions, but to gain confidence in choices they were already inclined to make. P2 and P4 sought confirmation for preliminary decisions across various situations, including outfit choices, housing options, and course schedules. Rather than seeking help with the initial decision-making process, they would present their analysis of options to the LLM, seeking reassurance to overcome their uncertainty. This pattern of seeking reassurance extended to impulse purchases, where P2 and P7 approached LLMs with specific expectations. As both noted, ”I just wanted to hear those words from someone” - revealing how users specifically sought out either permission or restraint from LLMs to support their predetermined decisions.
@@ 3.2.2. Finding the ”Right” Answer
Participants utilized LLMs to consider as many options and diverse sources of information to reach optimal decisions, and to check if their thought was the ”right” and ”optimal” answer. For instance, participants would validate their ideas before meetings or use LLMs to organize options for trip planning and gift buying. They believed LLMs would consider comprehensive contextual information to provide the ”right” answer. This aligned with participants’ views of ideal decision-making - to consider as many options as possible (P1, P4, P5, P6). Participants leveraged LLMs to strengthen this approach by having the models organize different options, though some (P5) expressed concern that their decision-making might be confined to LLM-suggested options, potentially missing other alternatives.
@@ 3.2.3. Optimizing Cognitive Resources through Task Delegation
Participants completely delegated tasks they saw as not worth their mental effort to LLMs, seeking to free themselves from such responsibilities while preserving their cognitive energy for more valuable activities. The definition of ”lower priority” varied among users in notably subjective ways. P6, a graduate student, chose to delegate course assignments to LLMs while engaging deeply with research work, despite both being essential academic responsibilities. P7 outsourced travel planning but enthusiastically handled laptop selection personally, citing a personal interest in electronics despite both tasks being similar consumer decisions. P3, another graduate student, delegated lab-related social interactions while actively engaging in research ideation, showing how even within the same professional context, participants made highly personal choices about which tasks to delegate. In these cases, participants accepted and implemented LLM suggestions, without seeking additional information or engaging in further analysis.
@@
3.3. RQ3: How do users conceptualize and evaluate their relationship with LLMs?
@@ 4. Discussion
Our study reveals how heavy users integrate LLMs into their daily tasks through distinct patterns. While delegating a broad range of decisions raised potential concerns about over-reliance and diminished critical thinking, our findings also highlight a nuanced form of human-AI collaboration where users and LLMs develop complementary relationships. Participants showed diverse usage strategies, treating LLMs as an emerging problem-solving tool and developing sophisticated prompting techniques. Most notably, participants frequently sought LLM guidance on social appropriateness and interpersonal situations. Although some users expressed concerns about potential skill degradation and a sense of unease, LLM consultations often led to a more thorough consideration of social factors and an enhanced understanding of different perspectives.
@@
This raises questions for future research on redefining how we conceptualize and measure over-reliance on LLMs. Current metrics typically assess over-reliance through simplified quantitative measures in controlled settings, primarily focusing on users’ acceptance rates of LLM outputs  (Bo et al., 2024; Kim et al., 2024). However, our findings reveal more complex patterns of engagement. Participants did not blindly adopt LLM outputs, even in cases where the eventually accepted them. Instead, participants demonstrated thoughtful delegation strategies, using LLMs to validate existing decisions, automate routine tasks, or navigate unfamiliar situations. The critical concern was not users’ acceptance of LLM outputs, but rather instances where users adopted LLM reasoning without exploring alternative perspectives. Future research should expand the definition of over-reliance beyond simple acceptance rates to examine how users critically engage with alternative lines of reasoning.
@@
Another key direction for future research involves capturing diverse user contexts. Our participants valued the ability of LLMs to extract necessary contextual information when not initially provided. They appreciated that they could receive meaningful responses without extensively explaining background information, even for context-heavy topics like relationship advice. Future research should explore ways to incorporate multimodal inputs beyond text-based interactions, allowing users to convey context through various channels. Additionally, LLMs’ ability to elicit implicit user intentions without explicit prompting is crucial, as demonstrated by recent advances in reasoning-focused LLM architectures that can proactively identify and address underlying user needs.
@@
The development of active usage patterns with LLMs appeared more prominent among younger users who had less experience managing tasks without these systems. Participants with extensive pre-LLM experience maintained clearer boundaries and showed greater awareness of system limitations. In contrast, users with less experience with LLMs demonstrated fewer reservations, viewing LLM interaction itself as a skill and actively developing their prompting strategies. Conducting design studies focused younger generations, to better understand and support these emerging interaction patterns represents a crucial direction for future research.

### v2

@@ Abstract.
copyright: noneconference: Tools for Thought: Research and Design for Understanding, Protecting, and Augmenting Human Cognition with Generative AI on CHI 2025 Workshop; April 26,2025; Yokohama, JAPANisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Recent studies have attempted to measure the impact of LLM use through quantitative metrics such as task performance and decision accuracy (Kim et al., 2025). However, these measurement-based approaches cannot fully capture how people delegate everyday decisions to LLMs or the resulting meta-cognitive effects. Furthermore, everyday decisions encompass a broad spectrum of choices, from routine task management to social interaction planning (Eigner and Händler, 2024; Dhami and Thomson, 2012), making them challenging to examine through purely quantitative and task-specific approaches.
@@
  • RQ2: What underlying needs do heavy LLM users seek to fulfill through LLM assistance?
  • RQ3: How do heavy LLM users conceptualize and evaluate their relationship with LLMs?
@@ 3.1.3. Case 3: Interpersonal Guidance through LLMs
3.2. RQ2: What underlying needs do heavy LLM users seek to fulfill through LLM assistance?
@@ 3.2.1. Gaining Self-Confidence through Validation
Participants often consulted LLMs not to make decisions, but to gain confidence in choices they were already inclined to make. P2 and P4 sought confirmation for preliminary decisions across various situations, including outfit choices, housing options, and course schedules. Rather than seeking help with the initial decision-making process, they would present their analysis of options to the LLM, seeking reassurance to overcome their uncertainty. This pattern of seeking reassurance extended to impulse purchases, where P2 and P7 approached LLMs with specific expectations. As both noted, ”I just wanted to hear those words from someone”—revealing how users specifically sought out either permission or restraint from LLMs to support their predetermined decisions.
@@ 3.2.2. Finding the ”Right” Answer
Participants utilized LLMs to consider as many options and diverse sources of information to reach optimal decisions, and to check if their thoughts were the ”right” and ”optimal” answer. For instance, participants would validate their ideas before meetings or use LLMs to organize options for trip planning and gift buying. They believed LLMs would consider comprehensive contextual information to provide the ”right” answer. This aligned with participants’ views of ideal decision-making - to consider as many options as possible (P1, P4, P5, P6). Participants leveraged LLMs to strengthen this approach by having the models organize different options, though some (P5) expressed concern that their decision-making might be confined to LLM-suggested options, potentially missing other alternatives.
@@ 3.2.3. Optimizing Cognitive Resources through Task Delegation
Participants completely delegated tasks they saw as not worth their mental effort to LLMs, seeking to free themselves from such responsibilities while preserving their cognitive energy for more valuable activities. The definition of ”lower priority” varied among users in notably subjective ways. P6, a graduate student, chose to delegate course assignments to LLMs while engaging deeply with research work, despite both being essential academic responsibilities. P7 outsourced travel planning but enthusiastically handled laptop selection personally, citing a personal interest in electronics despite both tasks being similar to consumer decisions. P3, another graduate student, delegated lab-related social interactions while actively engaging in research ideation, showing how even within the same professional context, participants made highly personal choices about which tasks to delegate. In these cases, participants accepted and implemented LLM suggestions without seeking additional information or engaging in further analysis.
@@
3.3. RQ3: How do heavy LLM users conceptualize and evaluate their relationship with LLMs?
@@ 4. Discussion
Our study reveals how heavy users integrate LLMs into their daily tasks through distinct patterns. Rather than simple tool usage, participants demonstrated sophisticated cognitive offloading strategies that transformed their decision-making processes. In our study, we observed participants delegating social and interpersonal reasoning to LLMs, suggesting ways users might leverage AI collaboration to support their social cognition processes.
Participants’ mental models of LLMs directly influenced their cognitive strategies—those viewing LLMs as rational entities engaged in cognitive complementarity by leveraging LLM capabilities where they perceived personal limitations, while those viewing LLMs as average decision-makers used cognitive benchmarking, establishing baseline standards while reserving higher-order tasks for themselves.
@@
This raises questions for future research on redefining how we conceptualize and measure over-reliance on LLMs. Current metrics typically assess over-reliance through simplified quantitative measures in controlled settings, primarily focusing on users’ acceptance rates of LLM outputs  (Bo et al., 2024; Kim et al., 2024). However, our findings reveal more complex patterns of engagement. Participants did not blindly adopt LLM outputs, even in cases where they eventually accepted them. Instead, participants demonstrated thoughtful delegation strategies, using LLMs to validate existing decisions, automate routine tasks, or navigate unfamiliar situations. The critical concern was not users’ acceptance of LLM outputs, but rather instances where users adopted LLM reasoning without exploring alternative perspectives. Future research should expand the definition of over-reliance beyond simple acceptance rates to examine how users critically engage with alternative lines of reasoning.
@@
Another key direction for future research involves capturing diverse user contexts. Our participants valued the ability of LLMs to extract necessary contextual information when not initially provided. They appreciated that they could receive meaningful responses without extensively explaining background information, even for context-heavy topics like relationship advice. Future research should explore ways to incorporate multi-modal inputs beyond text-based interactions, allowing users to convey context through various channels. Additionally, LLMs’ ability to elicit implicit user intentions without explicit prompting is crucial, as demonstrated by recent advances in reasoning-focused LLM architectures that can proactively identify and address underlying user needs.
@@
The development of active usage patterns with LLMs appeared more prominent among younger users who had less experience managing tasks without these systems. Participants with extensive pre-LLM experience maintained clearer boundaries and showed greater awareness of system limitations. In contrast, users with less experience with LLMs demonstrated fewer reservations, viewing LLM interaction itself as a skill and actively developing their prompting strategies. Conducting design studies focused on younger generations, to better understand and support these emerging interaction patterns represents a crucial direction for future research.